OpenStudy (anonymous):

A gas station stores its gasoline in a tank underground. The tank is a cylinder lying horizontally on its side. The radius is 4 ft, the length is 11 ft, and the top of the tank is 10 feet under the ground. The density of gasoline is 42 lb/ ft^3 Consider a slice of gasoline that is Δy ft thick and located y ft above the center of the cylinder. Write an expression giving the work required to pump the slice out

OpenStudy (anonymous):

\[ Work = Force \times Distance_{to\ the\ top} \]We let for be equal to the gravitational force because that is the minimum force needed to suck the gasoline out.\[ Force = Mass \times Acceleration_{gravity} \]Still need mass...\[ Mass = Volume \times Density \]They gave us density, but we need volume. Remember that we are trying to find the volume of a cylinder.\[ Volume = \pi \times Radius^2 \times Height \]Putting it all together, we get: \[ Work = \pi \times Radius^2\times Height\times Density\times Acceleration_{gravity} \times Distance_{to\ the\ top} \]